# Robustly learning the Hamiltonian dynamics of a superconducting quantum processor

### v2

@@ Methods
We use the Sycamore quantum processor composed of quantum systems arranged in a two-dimensional array. This processor consists of gmon qubits (transmons with tunable coupling) with frequencies ranging from 5 to 7 GHz. These frequencies are chosen to mitigate a variety of error mechanisms such as two-level defects. Our coupler design allows us to quickly tune the qubit–qubit coupling from 0 to 40+ MHz. The chip is connected to a superconducting circuit board and cooled down to below 20 mK in a dilution refrigerator. Each qubit has a microwave control line used to drive an excitation and a flux control line to tune the frequency. The processor is connected through filters to room-temperature electronics that synthesize the control signals. We execute single-qubit gates by driving 25 ns microwave pulses resonant with the qubit transition frequency.
@@
The qubits are connected to a resonator that is used to read out the state of the qubit. The state of all qubits can be read simultaneously by using a frequency-multiplexing. Initial device calibration is performed using ‘Optimus’ Kelly et al. (2018) where calibration experiments are represented as nodes in a graph.
@@
$y_{m,n}[l]:=\langle a_{m}(t_{l})\rangle_{\psi_{n}}=\frac{1}{2}(M\cdot\exp(-\mathrm{i}t_{l}h)\cdot S)_{m,n},$ (6)
@@
where we have dropped $S$ and $M$ for the time being. This decomposition suggests a simple procedure to identify the Hamiltonian using Fourier data analysis. From the matrix-valued time series data $y[l]$ (7), we identify the Hamiltonian coefficient matrix $h$ in two steps. First, we determine the eigenfrequencies of $h$. Second, we identify the eigenbasis of $h$. To achieve those identification tasks with the largest possible robustness to error, it is key to exploit all available structure at hand.
@@
In order to robustly estimate the spectrum, we exploit that the signal is sparse in Fourier space. This structure allows us to substantially denoise the signal and achieve super-resolution beyond the Nyquist limit Candès and Fernandez-Granda (2013, 2014). A candidate algorithm for this task, suitable for scalar time-series, is the ESPRIT algorithm, which comes with rigorous recovery guarantees Fannjiang (2016); Li et al. (2020b). To extract the Hamiltonian spectrum from the matrix time-series $y[l]$, we apply ESPRIT to the trace of the data series (for $S=M=\mathbbm{1}$)
@@
To overcome this issue and obtain a truly scalable learning procedure applicable to degenerate spectra, we develop a new algorithm coined tensorESPRIT, which extends the ideas of ESPRIT to the case of a matrix time-series using tensor network techniques. Using tensorESPRIT also improves the robustness of frequency estimation to SPAM errors. For practical Hamiltonians, tensorESPRIT becomes necessary for systems with $N\gtrsim 12$.
@@
To identify the eigenspaces of the Hamiltonian, we use the eigenfrequencies found in Step 1 to fix the oscillating part of the dynamics in Eq. 7. What remains is the problem of finding the eigenspaces $\left.\left|{v_{k}}\right.\right\rangle\!\!\left.\left\langle{v_{k}}\right.\right|$ from the data. This problem is a non-convex inverse quadratic problem, subject to orthogonality of the eigenspaces, as well as the constraint that the resulting Hamiltonian matrix respects the connectivity of the superconducting architecture. Formally, we denote the a priori known support set of the Hamiltonian matrix as $\Omega$, so that we can write the support constraint as $h_{\bar{\Omega}}=0$, where $\bar{\Omega}$ denotes the complement of $\Omega$ and subscripting a matrix with a support set restricts the matrix to this set. We can cast this problem into the form of a least-squares optimization problem
@@
equipped with non-convex constraints enforcing orthogonality, and the quadratic constraint restricting the support. In order to approximately enforce the support constraint, we make use of regularization Bühlmann and Geer (2011). It turns out that this can be best achieved by adding a term (Hangleiter et al., 2020, App. A)
@@
to the objective function (9), where $\mu>0$ is a parameter weighting the violation of the support constraint. We then solve the resulting minimization problem by using a conjugate gradient descent on the manifold of the orthogonal group Edelman et al. (1998); Abrudan et al. (2009), see also the recent work Luchnikov et al. (2020, 2021); Roth et al. (2023) for the use of geometric optimization for quantum characterization.
@@
The caveat of this approach is that the noise that affected $y[l_{0}]$ is now present in every entry of the new data series $y^{(l_{0})}$. We can remedy this effect by using several time points $y[l_{0}]$ for the inversion. To this end, we compute the concatenation of data series for different choices of $l_{0}$, e.g., for every $s$ data points $0,s,2s,\ldots,\lfloor L/s\rfloor s$ giving rise to new data $y_{\text{total, s}}=(y^{(0)},y^{(s)},y^{(2s)},\ldots,y^{(\lfloor L/s\rfloor s)})\in\mathbb{C}^{\lfloor L/s\rfloor L}$. If the data suffers from drift errors, it is also beneficial to restrict each data series $y^{(l_{0})}$ to entries $y^{(l_{0})}[\kappa]$ with $\kappa\in[l_{0}-w,l_{0}+w]$, i.e., the entries in a window of size $w$ around $l_{0}$. In practice, we use $s=1$ and $w=50$ for the reconstructions on Sycamore #1, and $s=1,w=L$ for those on Sycamore #2.
As we argue below, the final map $M$ is nearly diagonal here. Hence, we can use $y_{\text{total, s}}$ from Eq. (9) and it is justified to apply the support constraint in the eigenspace reconstruction step. However, the eigenspace reconstruction will suffer from systematic errors due to the final map, even in the case when it is nearly diagonal. Below, we explain a method to partially remove this error.
@@
where $\tilde{h}=\sum_{k}\lambda_{k}\left.\left|{v_{k}}\right.\right\rangle\!\!\left.\left\langle{v_{k}}\right.\right|$ and $\{\left.\left|{v_{k}}\right.\right\rangle\}$ is the eigenbasis obtained by solving the problem (9), and use $\hat{D}_{M}$ as an estimate of $M$. We can now obtain a tomographic estimate of the initial map through
$\hat{S}:=\frac{2}{L+1}\sum_{l=0}^{L}\exp[\mathrm{i}t_{l}\hat{h}]\hat{D}_{M}y[l]\,.$ (13)
@@
We benchmark our identification algorithm against more direct approaches in numerical simulations including models for statistical and systematic errors in the SM VI. We find that, indeed, already for small system sizes, the regularized manifold optimization algorithm developed here features an improved robustness against state preparation and measurement errors compared to (post-projected) linear inversion. For intermediate system sizes ($N>10$), exploiting structure in the recovery algorithm then becomes an imperative. In particular, for larger system sizes the eigenspectrum of the Hamiltonian becomes unavoidably narrower spaced. We find that on instances of the Harper Hamiltonian studied here linear inversion approaches cannot be applied at all for $N>20$. Regularized conjugate gradient decent in contrast yields good recovery performance even for larger systems. The same limitations apply to a direct Fourier analysis of the cumulative time series data using ESPRIT, as described above. For different families of Hamiltonians, we find that above a system size of $N\approx 20$ tensorESPRIT still consistently recovers the frequency spectrum, while the ESPRIT algorithm fails to do so.
@@ Acknowledgements.
D.H. and I.R. conceived of the Hamiltonian identification algorithm. J.F. conceived of the tensorESPRIT algorithm. D.H., I.R. and J.F. analyzed the experimental data and benchmarked the identification algorithm. P.R. took the experimental data. D.H. and I.R. wrote the initial manuscript. All authors contributed to discussions and writing the final manuscript.

### v3

@@ Methods
We use the Sycamore quantum processor composed of quantum systems arranged in a two-dimensional array. This processor consists of gmon qubits (transmons with tunable coupling) with frequencies ranging from $5$ to $7$ GHz. These frequencies are chosen to mitigate a variety of error mechanisms such as two-level defects. Our coupler design allows us to quickly tune the qubit-qubit coupling from 00 to $40+$ MHz. The chip is connected to a superconducting circuit board and cooled down to below $20$ mK in a dilution refrigerator. The median values of the $T_{1}$ and $T_{2}$ times of the qubits are $T_{1}=16.1$ µs, $T_{2}=5.3$ µs (Ramsey interferometry) and $T_{2}=17.8$ µs (after CPMG dynamical decoupling). Each qubit has a microwave control line used to drive an excitation and a flux control line to tune the frequency. The processor is connected through filters to room-temperature electronics that synthesize the control signals. We execute single-qubit gates by driving 25 ns microwave pulses resonant with the qubit transition frequency, resulting in single-qubit gate fieldity of 99.8% as measured via randomized benchmarking.
Ramping pulses.
The pulses used in the experiment are pre-distorted in order to compensate for filters on the control lines. In order to calibrate this distortion, we send rectangular pulses to the qubits and monitor the frequency change of the qubits. This allows us to know the response of the microwave lines at the qubits (i.e., the deviation from a rectangle) and compensate for distortions. The ramp time can be as fast as 2 to 3 ns and the distortions take the form of overshoot and undershoots with a long response time of  100ns. After compensating for the distortions, the qubit frequency remains fixed.
@@
The qubits are connected to a resonator that is used to read out the state of the qubit. The state of all qubits can be read simultaneously by using a frequency-multiplexing. Initial device calibration is performed using ‘Optimus’ kelly_physical_2018 where calibration experiments are represented as nodes in a graph.
@@
$y_{m,n}[l]\coloneqq\langle a_{m}(t_{l})\rangle_{\psi_{n}}=\frac{1}{2}(M\cdot\exp(-\mathrm{i}t_{l}h)\cdot S)_{m,n},$ (6)
@@
where we have dropped $S$ and $M$ for the time being. This decomposition suggests a simple procedure to identify the Hamiltonian using Fourier data analysis. From the matrix-valued time series data $y[l]$ (S2), we identify the Hamiltonian coefficient matrix $h$ in two steps. First, we determine the eigenfrequencies of $h$. Second, we identify the eigenbasis of $h$. To achieve those identification tasks with the largest possible robustness to error, it is key to exploit all available structure at hand.
@@
In order to robustly estimate the spectrum, we exploit that the signal is sparse in Fourier space. This structure allows us to substantially denoise the signal and achieve super-resolution beyond the Nyquist limit CandesFernandez-Granda:2013; CandesFernandez-Granda:2014. A candidate algorithm for this task, suitable for scalar time-series, is the ESPRIT algorithm, which comes with rigorous recovery guarantees Fannjiang:2016:ESPRIT; LiLiaoFannjiang:2019:Super-resolution. To extract the Hamiltonian spectrum from the matrix time-series $y[l]$, we apply ESPRIT to the trace of the data series (for $S=M=\mathbbm{1}$)
@@
To overcome this issue and obtain a truly scalable learning procedure applicable to degenerate spectra, we develop a new algorithm coined tensorESPRIT. TensorESPRIT extends the ideas of ESPRIT to the case of matrix-valued time series using tensor network techniques. Using tensorESPRIT also improves the robustness of frequency estimation to SPAM errors. For practical Hamiltonians, tensorESPRIT becomes necessary for systems with $N\gtrsim 12$; as we find in numerical simulations summarized in Section .4 and detail in Section IV.B of the SM.
@@
To identify the eigenspaces of the Hamiltonian, we use the eigenfrequencies found in Step 1 to fix the oscillating part of the dynamics in Eq. S2. What remains is the problem of finding the eigenspaces $\left.\left|{v_{k}}\right.\right\rangle\!\!\left.\left\langle{v_{k}}\right.\right|$ from the data. This problem is a non-convex inverse quadratic problem, subject to orthogonality of the eigenspaces, as well as the constraint that the resulting Hamiltonian matrix respects the connectivity of the superconducting architecture. Formally, we denote the a priori known support set of the Hamiltonian matrix as $\Omega$, so that we can write the support constraint as $h_{\bar{\Omega}}=0$, where $\bar{\Omega}$ denotes the complement of $\Omega$ and subscripting a matrix with a support set restricts the matrix to this set. We can cast this problem into the form of a least-squares optimization problem
@@
equipped with non-convex constraints enforcing orthogonality, and the quadratic constraint restricting the support. In order to approximately enforce the support constraint, we make use of regularization Regularization. It turns out that this can be best achieved by adding a term (HangleiterEtAl:2020, App. A)
@@
to the objective function (S24), where $\mu>0$ is a parameter weighting the violation of the support constraint. We then solve the resulting minimization problem by using a conjugate gradient descent on the manifold of the orthogonal group EdelmanAriasSmith:1998; abrudan_conjugate_2009, see also the recent work luchnikov2020qgopt; luchnikov2020riemannian; RothEtAl:2020 for the use of geometric optimization for quantum characterization.
@@
The caveat of this approach is that the shot noise that affected the single time point $y[l_{0}]$ can lead to correlated errors in every entry of the new data series $y^{(l_{0})}$.
We can reduce the error induced by these correlations by effectively averaging over ‘corrected’ data series $y^{(l_{0})}$ with different $l_{0}$. To this end, we compute the concatenation of data series for different choices of $l_{0}$, e.g., for every $s$ data points $0,s,2s,\ldots,\lfloor L/s\rfloor s$ giving rise to new data $y_{\text{total, s}}=(y^{(0)},y^{(s)},y^{(2s)},\ldots,y^{(\lfloor L/s\rfloor s)})\in\mathbb{C}^{\lfloor L/s\rfloor L}$. If the data suffers from drift errors, it is also beneficial to restrict each data series $y^{(l_{0})}$ to entries $y^{(l_{0})}[\kappa]$ with $\kappa\in[l_{0}-w,l_{0}+w]$, i.e., the entries in a window of size $w$ around $l_{0}$. In practice, we use $s=1$ and $w=50$ for the reconstructions on Sycamore #1, and $s=1,w=L$ for those on Sycamore #2.
As we argue below, the final map $M$ is nearly diagonal here. Hence, we can use $y_{\text{total, s}}$ from Eq. (S24) and it is justified to apply the support constraint in the eigenspace reconstruction step. However, the eigenspace reconstruction will suffer from systematic errors due to the final map, even in the case when it is nearly diagonal. Below, we explain a method to partially remove this error.
@@
where $\tilde{h}=\sum_{k}\lambda_{k}\left.\left|{v_{k}}\right.\right\rangle\!\!\left.\left\langle{v_{k}}\right.\right|$ and $\{\left.\left|{v_{k}}\right.\right\rangle\}$ is the eigenbasis obtained by solving the problem (S24), and use $\hat{D}_{M}$ as an estimate of $M$. We can now obtain a tomographic estimate of the initial map through
$\hat{S}\coloneqq\frac{2}{L+1}\sum_{l=0}^{L}\exp[\mathrm{i}t_{l}\hat{h}]\hat{D}_{M}y[l]\,.$ (13)
@@
We benchmark our identification algorithm against more direct approaches in numerical simulations including models for statistical and systematic errors in the SM VI. We find that, indeed, already for small system sizes, the regularized manifold optimization algorithm developed here features an improved robustness against state preparation and measurement errors compared to (post-projected) linear inversion. For intermediate system sizes ($N>10$), exploiting structure in the recovery algorithm then becomes an imperative. In particular, for larger system sizes the eigenspectrum of the Hamiltonian becomes unavoidably narrower spaced, leading to (near-)degeneracies. We find that on instances of the Harper Hamiltonian studied here linear inversion approaches cannot be applied at all for $N>20$. Regularized conjugate gradient decent in contrast yields good recovery performance even for larger systems. The same limitations apply to a direct Fourier analysis of the cumulative time series data using ESPRIT, as described above. For different families of Hamiltonians, we find that above a system size of $N\approx 20$ tensorESPRIT still consistently recovers the frequency spectrum, while the ESPRIT algorithm fails to do so.
@@ Acknowledgements.
D.H. and I.R. conceived of the Hamiltonian identification algorithm. J.F. conceived of the tensorESPRIT algorithm. D.H., I.R. and J.F. analyzed the experimental data and benchmarked the identification algorithm. P.R. took the experimental data. D.H. and I.R. wrote the initial manuscript. D.H., I.R., J.F., J.E. and P.R. contributed to discussions and writing the final manuscript.